OpenStudy (anonymous):

Alicia bikes 25% of a 100-mi trip on the first day. She bikes 1/3 of the remaining distance on the second day. On the third day, she bikes 40% of the remaining distance. How many miles are left in Alicia’s trip

OpenStudy (anonymous):

100*(25/100)+75/3+50*(40/100)=25+25+20=70

OpenStudy (anonymous):

they refer to the reaming distances so it is 30 u can,t just multiply it

OpenStudy (anonymous):

i hope there is a mistake with u r answer Alicia bikes 25% of a 100-mi trip on the first day ie she bikes 25 then the 1/3 of reaming distance is covered 2nd day ie 75/3=25 then 40% of last fifty meter =20 so she traveled 70 so renaming is 30